Solve the following equations:
3x+4(x+2)=11+7x

Answer by Boreal(15235)

3x+4(x+2)=11+7x
Distribute
3x+4x+8=11+7x
collect terms
7x+8=11+7x
subtract 7x from both sides
8=11 NOT TRUE. No solution